A father is 25 years older than his son. In 10 years, he will be twice as old as his son.
Find their present ages.

Answer by ikleyn(52788)

.

Let x be the age of the son; then the father's age is (x+25), according to the condition.


In 10 years, the son's age will be (x+10), while the father's age will be (x+25+10) = x+35.


The condition says that

    2*(x+10) = x+35.


Simplify and solve

    2x + 20 = x + 35

    x       = 35 - 20 = 15.


ANSWER.  The son is 15 years old; the father is 15+25 = 40 years old.

Solved.

-----------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.